Sherris Tan-O-Rama is a local tanning salon. The following information reflects its number of appointments and total costs for the first half of the year: Total Cost $5,000 6,000 6,200 5,300 5,450 5,230 Number of Month January February March April May June 250 300 500 225 150 175 Using the high-low method, calculate the total fixed cost per month and the variable cost per tanning appointment. (Round your Variable Cost per Unit answer to 2 decimal places and Fixed Cost answer to the nearest dollar amount. Variable Cost per Unit Fixed Cost
Answer #1

Variable cost per unit = (6200-5450)/(500-150) = 2.14 per unit

Fixed cost = Total cost-variable cost

                 = 5450-(150*2.14)

Fixed cost = 5129